How many milliliters of a 25% hydrogen peroxide solution and how many milliliters of a 9% hydrogen peroxide solution should be mixed to get 176 mL of a 18% hydrogen peroxide solution?
Answer by lwsshak3(11628)

How many milliliters of a 25% hydrogen peroxide solution and how many milliliters of a 9% hydrogen peroxide solution should be mixed to get 176 mL of a 18% hydrogen peroxide solution?
***
let x=amt of 25% solution to mix
176-x=amt of 9% solution to mix
..
25%x+9%(176-x)=18%*176
.25x+15.84-.09x=31.68
.16x=15.84
x=99
176-x=77
amt of 25% solution to mix=99 mm
amt of 9% solution to mix=77 mm